OpenStudy (anonymous):

There are 5 boys and 7 girls on the swimming team. The ratio of boys to girls on the tennis team is proportional to the ratio of boys to girls on the swimming team. Which could be the number of boys and girls on the tennis team? A. 6 boys and 8 girls B. 8 boys and 10 girls C.

OpenStudy (anonymous):

not those two for sure

OpenStudy (anonymous):

I don't see the answer either.. What was letter C's answer?

OpenStudy (anonymous):

Sorry Here There are 5 boys and 7 girls on the swimming team. The ratio of boys to girls on the tennis team is proportional to the ratio of boys to girls on the swimming team. Which could be the number of boys and girls on the tennis team? A 6 boys and 8 girls . 6 boys and 8 girls B. 8 boys and 10 girls C. 10 boys and 14 girls

OpenStudy (anonymous):

Ok I Know Yall Probably Confused So I Made Sure I Wrote It Down Right. These Are The Answer Choices A. 6 boys and 8 girls B. 8 boys and 10 girls C. 10 boys and 12 girls D. 10 boys and 14 girls

OpenStudy (anonymous):

find the fraction \[\frac{boys}{girls}\] that is equal to \[\frac{5}{7}\]

OpenStudy (anonymous):

for example, it is not C because \[\frac{10}{12}=\frac{5}{6}\neq \frac{5}{7}\]

OpenStudy (anonymous):

So It Would Be D?